OpenStudy (anonymous):

Let \(f_0(x)=e^x \) and \(f_{n+1} (x)=xf_n'(x) \) for \(n=0,1,2,... \) ........Find the value of\[\sum_{n=0}^{\infty} \frac{f_n(1)}{n!}\]

OpenStudy (dominusscholae):

e^2?

OpenStudy (anonymous):

that was close...nope...

OpenStudy (experimentx):

the coefficients are exp(1) exp(1) 2 exp(1) 5 exp(1) 15 exp(1) 52 exp(1) 203 exp(1) 877 exp(1) 4140 exp(1) 21147 exp(1) 115975 exp(1)

OpenStudy (anonymous):

\[e^x=\sum_{k=0}^{\infty} \frac{x^k}{k!}\]

OpenStudy (dominusscholae):

Here's my work so far.